OpenStudy (ny,ny):

Find the values of x that give relative extrema for the function f(x) = (x+2)^2(x-2).

OpenStudy (ny,ny):

I think I have to find the derivative first. But I don't know how to do that.

OpenStudy (jabberwock):

You will need to use the product rule. (First function)(Derivative of the second function) + (Derivative of first function)(Second function)

OpenStudy (ny,ny):

F'(x) = (x+1)^2(1)+(2x+2)(x-2) Like this?

OpenStudy (steve816):

You won't need the product rule if you expand the function.

OpenStudy (jabberwock):

Ah, true @steve816

OpenStudy (ny,ny):

Ok. Ill expand right now.

OpenStudy (ny,ny):

x^3-3x-2 ?

OpenStudy (jabberwock):

Your answer is correct @NY,NY except that the second term should be 2(x+2)(x-2)

OpenStudy (jabberwock):

*of your previous answer

OpenStudy (jabberwock):

\[(x+2)^2(x-2)=(x^2+4x+4)(x-2)\]

OpenStudy (jabberwock):

\[=x^3+4x^2+4x-2x^2-8x-8\]

OpenStudy (jabberwock):

\[=x^3+2x^2-6x-8\]

OpenStudy (jabberwock):

Does that look right, or did I make a mistake?

OpenStudy (jabberwock):

*By the way, both these methods should give you the same answer.

OpenStudy (ny,ny):

Oh i was so confused but I just realized i typed the question wrong F(x) is (x+1)^2(x-2) So sorry about that.

OpenStudy (jabberwock):

Ah, okay.

OpenStudy (jabberwock):

Okay, I got what you got

OpenStudy (jabberwock):

The nice thing about this is that you can take the derivative of each of the terms individually.

OpenStudy (ny,ny):

Yes You mean with x^3-3x-2 right?

OpenStudy (ny,ny):

Because we would take the derivative idividually now To become 3x^2-3 And x = 1, -1

OpenStudy (jabberwock):

Absolutely. :)

OpenStudy (ny,ny):

Thank you.

OpenStudy (jabberwock):

No worries. Good luck.